OpenStudy (anonymous):

5^x=4^x+1 becomes (x)ln5=(x+1)ln4 ....where do I go from here?

OpenStudy (anonymous):

(ln5/ln4)=(x+1/x)?

OpenStudy (anonymous):

(ln5/ln4)=(x+1/x)?

OpenStudy (anonymous):

distribute the right side

OpenStudy (anonymous):

remember that ln(4) is just a nuumber

OpenStudy (anonymous):

still having trouble. ln5/ln4 is 1.1609=(x+1/x)

OpenStudy (anonymous):

so you have, xln(5) = xln(4) + ln(4)

OpenStudy (anonymous):

x ln(5) - xln(4) = ln(4) x [ln(5) - ln(4)] = ln(4) ln(4) x = --------------- ln(5) - ln(4)

OpenStudy (anonymous):

that was just what I needed. Thanks a lot!!